OpenStudy (anonymous):

Use the distance formula. P(-2, 7) and Q(-6, 4)

OpenStudy (saifoo.khan):

When u know the formula, simply use it.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@saifoo.khan im trying to see if im correct, because my answer was \[\sqrt{-11}\]

OpenStudy (anonymous):

formula sqr root of [ (xsub2-xsub1)^2+(ysub2-ysub1)^2 ] same for 3 dimensions except you include the variable z sqr root of [ (xsub2-xsub1)^2+(ysub2-ysub1)^2 +(zsub2-zsub1)^2]

OpenStudy (anonymous):

i could be wrong, havent looked to see

OpenStudy (anonymous):

P(-2, 7) and Q(-6, 4) d=√((x^2-x^1 )+(y^2-y^1 ) ) =√([-6-(-2) ]+(4-7)) =√(-8+3) =√(-5) =5 kk i see what i did wrong now.

OpenStudy (anonymous):

over 4, down 3 pythagoras says it is \[d^2=3^2+4^2=25\implies d = 5\]